Weight 0.153N determine the specific weight, density and specific gravity of the juice express you answer in si unit


1
Expert's answer
2021-01-18T01:19:01-0500

specific weight of juice = 0.153 / 9.81 = 0.0156 kg

density = m/V = 0.0156 kg / 0.0146 x 10-3 m3 = 1 080 kg/m3

specific gravity = density x g = 1.08 x 9.81 = 10.6 kN/m3